# Deep learning for classifying dynamical states from time series via recurrence plots

### v1

@@ II.3 Recurrence Plots
Choosing an appropriate threshold $\epsilon$ is critical for constructing meaningful recurrence plots, as it determines which state pairs are considered recurrent and thereby influences the actual dynamical state of the system. An optimal threshold preserves significant features of the dynamics while excluding noise-induced or irrelevant recurrences [40]. If $\epsilon$ is too small, the recurrence plot may contain very few recurrence points, leading to the omission of important dynamical patterns. Conversely, if $\epsilon$ is too large, the plot becomes overly dense, which can obscure important structures such as diagonal lines that reflect periodic or quasi-periodic behavior.
@@
To select an appropriate threshold, common approaches include fixing $\epsilon$ as a percentage of the signal’s standard deviation [41], or adjusting it to maintain a fixed recurrence rate across datasets [21]. In our study, we set the threshold $\epsilon$ to achieve a fixed recurrence rate of $5\%$. To do this, we compute all pairwise distances and select the distance value at which $5\%$ of the distances are smaller than or equal to it, as the threshold $\epsilon$.
@@
Figure 2: Representative recurrence plots. These plots are obtained from the univariate time series corresponding to the following six classes: (a) Periodic (Lorenz 4D), (b) Quasiperiodic, (c) Chaotic (Lorenz 4D), (d) Hyperchaotic (Chen 4D), (e) Red Noise, and (f) White Noise.
II.4 Deep Learning Models for Classification
@@
We first implement a single-branch deep learning model based on the ResNet-50 architecture as a baseline to compare with our proposed dual-branch model. This baseline model utilizes a pre-trained ResNet-50 network [31], trained on the ImageNet dataset[42], to extract spatial features from input RPs. We apply a Global Average Pooling (GAP) layer to the final convolutional output of the ResNet-50 backbone. The resulting pooled features are passed through a fully connected layer with 256 neurons and ReLU activation to capture nonlinear relationships. To avoid overfitting, a dropout layer is applied with a rate of $0.5$. Finally, a softmax output layer produces a probability distribution over the target classes.
Building upon this, we implement a dual-branch deep learning model inspired by the DSRNet architecture [32]. As in the baseline, a pre trained ResNet-50 backbone (trained on ImageNet) is used to extract feature maps from the input RPs. These feature maps are then processed through two parallel branches. The standard branch mirrors the baseline model, applying a GAP layer to the final convolutional output of the ResNet-50 backbone. The directional branch begins by reducing the depth of the ResNet feature maps to a single channel using a $1\times 1$ convolution. Following the design of DSRNet, eight fixed $3\times 3$ directional convolution kernels are applied to highlight edge information along eight principal orientations: horizontal (0°), vertical (90°), and diagonals (45°, 135°, 180°, 225°, 270°, 315°). These directionally filtered maps are further processed by a learnable $3\times 3$ convolution layer, followed by GAP.
The outputs of the standard and directional branches are concatenated and passed through a fully connected layer with $256$ neurons and ReLU activation. A Dropout layer (rate = $0.5$) is again applied for regularization, and the final softmax layer maps the combined representation to the class probabilities.
III Results
The goal of this work is to accurately identify and classify distinct dynamical behaviors and noisy patterns using our developed machine learning model trained on RPs. We generate a total of 1800 RPs per class from various dynamical systems and noises, as shown in Table 1. Each deterministic class includes 800 RPs from the respective system and 800 noise-augmented variants derived from the same system, as described in Section II.1. The complete dataset contains six distinct classes. Representative examples of RPs from each class are presented in Figure 2. The dataset is partitioned into training and testing sets using an 80:20 split, ensuring that the testing data remains completely unseen during the training process. All images are resized to $800\times 800$ pixels and converted to float16 precision to optimize memory usage. Model training is performed using the Adam optimizer with a low learning rate $10^{-7}$ to ensure stable convergence. To prevent overfitting, early stopping is applied with a patience parameter of $2$, monitored on the test error. The model is trained for a maximum of $100$ epochs with a batch size of $8$.
@@ III Results
To enhance prediction accuracy, we apply a test-time augmentation strategy in which each input image is rotated in multiple orientations and cropped by up to $10\%$, while preserving the original aspect ratio. Predictions are generated for all augmented variants of the same image, and the class with the highest average confidence across these predictions is selected as the final output. This ensemble-like approach improves robustness to minor variations in recurrence plot appearance and mitigates overfitting to specific orientations or features.
@@
To assess the performance of the classification scheme, we compute per-class accuracy in addition to confusion matrices and standard classification metrics, as follows: Given a confusion matrix $\mathbf{C}\in\mathbb{N}^{K\times K}$, where $C_{i,j}$ denotes the number of samples whose true class is $i$ and predicted class is $j$, and $K$ denotes the number of classes, the accuracy per class for class $i$ is defined as:
@@
where, $C_{i,i}$ is the number of correctly predicted samples for the class $i$, $\sum_{j=1}^{K}C_{i,j}$ is the total number of samples belonging to the class $i$. This metric reflects how well the model performs individually for each class.
III.1 Classification of simulated time series data
To evaluate the performance of both trained machine learning models, ResNet-50 and our proposed dual-branch model, we first validate them on simulated time series data drawn from the same distribution as the training set, generated using the systems listed in Table 1. Both models demonstrate comparable performance on this simulated dataset and therefore, we present only the results obtained from our dual-branch model.
The confusion matrix in Figure 3 provides a representative overview of classification performance across all six classes. The diagonal elements indicate correct classifications, while off-diagonal elements represent misclassifications. Figure 3 shows that all predictions lie on the diagonal, indicating perfect classification of the test dataset. The evaluation metrics of the model, including precision, recall, F1 score, and per-class accuracy, all indicate 100% performance, further confirming that the model achieves an overall accuracy of $100\%$.
Given the perfect accuracy, we check the possibility of model’s overfitting. Overfitting typically occurs when a model memorizes the training data and fails to generalize, often indicated by a divergence between train and test accuracies during training. However, as shown in Figure 4, both train and test accuracies remain consistently high and closely aligned throughout the training process, suggesting that overfitting is unlikely in our case. This robustness may be attributed to the similarity between the train and test datasets, as well as the use of early stopping, which terminates training before the onset of overfitting.
Figure 3: Confusion matrix. The matrix illustrates the classification performance of the dual-branch model on simulated test datasets. The diagonal cells show how many recurrence plots are correctly classified, while the off-diagonal cells show how many are misclassified into other classes.
Figure 4: Comparison of accuracy and error rates between the training and testing phases for the dual-branch model Our dual-branch model exhibits similar trends and convergence in accuracy and error for both training and testing time series data, indicating that the model does not overfit.
To further validate the model using simulated datasets, we evaluate its performance on data generated from the Lorenz 3D[43] and Rössler 3D[44] systems, which are not included during the training phase. For the Lorenz 3D system, we generate time series from both chaotic and periodic regimes and construct the corresponding recurrence plots. These plots are accurately classified as chaotic and periodic, respectively, by both the ResNet-50 model and the proposed dual-branch model. Similarly, for the Rössler 3D system, we generate recurrence plots from chaotic and periodic regimes, which are also correctly identified by both models.
III.2 Classification of empirical time series
@@
To further evaluate the generalizability of our trained deep learning model, we test it on experimental time series data. As a first example, we use the time series data obtained from the Chua circuit . We generate and pass the recurrence plots of different time series lengths to the models in order to test their performance on varying pattern scales within the recurrence plots. The cropping and the corresponding recurrence plots are shown in Figure 5. Both ResNet-50 and the proposed dual-branch model correctly classify the Chua circuit signal as chaotic, consistent with the well-established chaotic dynamics of the Chua system [5, 45].
We further validate the model using the light curve data from three variable stars, AC Her, SX Her, and Chi Cygni, sourced from the AAVSO database (https://www.aavso.org/data-download). Figure 6 presents the light curves and their corresponding recurrence plots. The light curves for AC Her and SX Her span 14 years, beginning on 20 September 1980, while the Chi Cygni light curve covers a 5-year period starting on 20 September 2019. To ensure consistent sampling and reduce observational noise, we smooth the AC Her and SX Her light curves using a 5-day rolling average, and the Chi Cygni light curve using a 20-day rolling average. All light curves are then interpolated using cubic splines with a smoothing parameter of $s=0.065$.
The light curve of Chi Cygni is correctly classified as periodic by both the ResNet-50 and dual-branch models, in agreement with previous studies [46]. However, the ResNet-50 model misclassifies AC Her as white noise and SX Her as hyperchaotic, contradicting earlier findings that identify both stars as exhibiting chaotic dynamics [47, 48]. In contrast, our proposed dual-branch model correctly classifies both AC Her and SX Her as chaotic, in alignment with prior literature. Hence, we conclude that the proposed dual-branch model provides a more reliable framework than the ResNet-50 model for accurately identifying the underlying dynamics of empirical time series.
@@
Figure 5: Experimental data from Chua’s circuit. (a) Time series obtained experimentally from Chua’s circuit. The red line indicates the cropped portion of the timeseries (b) Corresponding RP and the red box highlights the region in the RP corresponding to only the cropped portion shown in (a).
@@
IV Conclusion
In this work, we introduce a new deep learning model that accurately predicts the inherent dynamical nature of both synthetic and empirical time series data. The central question we address is one of broad relevance: Can we reliably infer the underlying dynamics of real-world time series data? To answer this, we integrate recurrence plots, extracted via phase-space reconstruction using Takens’ embedding theorem, with deep learning models that learn to identify complex dynamical structures directly from the RPs. This approach demonstrates high accuracy and strong generalizability, making it a scalable and robust solution for dynamical states classification.
@@ IV Conclusion
We implement a customized dual-branch deep learning model that takes RP images as inputs. These RPs are generated from nonlinear time series obtained through simulations of various dynamical systems and noise data. Our model can classify these time series into six distinct classes: periodic, quasi-periodic, chaotic, hyperchaotic, white noise, and red noise. The architecture of our model builds on a ResNet-50 backbone and incorporates two parallel branches: one that captures global spatial features and another that extracts directional and edge-based patterns, enhancing the model’s classification performance over the standard ResNet-50.
Table 2: Comparison of the performance of ResNet-50 and the new dual-branch model in the classification of unseen simulated and experimental time series data
Time Series ResNet-50 Dual-Branch Model True Class
Simulated Time Series Data
Lorenz 3D Chaotic Chaotic Chaotic [43]
Periodic Periodic Periodic [43]
Rossler 3D Chaotic Chaotic Chaotic [44]
Periodic Periodic Periodic [44]
Experimental/Observational Time Series Data
Chua Circuit Chaotic Chaotic Chaotic [5]
Chi Cygni Periodic Periodic Periodic [46]
AC Her White Noise Chaotic Chaotic [48]
SX Her Hyperchaotic Chaotic Chaotic [47]
In summary, our framework demonstrates robust classification performance across synthetic, experimental, and observational datasets. Notably, the trained dual-branch model accurately classifies unseen simulated time series from the Lorenz 3D and Rössler 3D systems, as well as experimental data from a Chua’s circuit and astronomical light curves from variable stars AC Her, SX Her, and Chi Cygni, as shown in Table 2. This demonstrates the model’s applicability to real-world data and reinforces the potential of using recurrence based deep learning models for determining the inherent dynamics of the system. We observe that our dual-branch model outperforms than other convolutional neural networks models that have been used to classify the dynamical behavior using RPs as inputs.
@@
We also compared with traditional approaches that rely on recurrence quantification measures and recurrence network metrics [24], our method offers a significantly more efficient alternative. It eliminates the need for manual feature engineering by learning discriminative features directly from RP images. Our results show that high classification accuracy can be achieved without computing various recurrence measures.
@@
Looking ahead, several directions could further extend the capabilities of our model. A promising approach involves detecting dynamical transitions within time series by segmenting recurrence plots along the line of identity (LOI) using our model on transition-rich data to identify these substructures, similar to time-dependent RQA techniques [40]. Another enhancement includes combining our approach with model-based generative frameworks, such as delay embedding and deep delay autoencoders [49], to synthesize training data from partial observations of unknown systems. This could reduce dependence on fully simulated datasets and improve performance under limited data conditions.
In conclusion, this work establishes a new deep learning model that takes recurrence plots as images and offers a powerful, efficient, and generalizable framework for classifying dynamical states. It opens new possibilities for analyzing complex time series in diverse fields such as physics, neuroscience, astrophysics, and beyond.
V Acknowledgments

### v2

@@ 2.3 Recurrence Plots
The recurrence plot is then derived by adding dots corresponding to the nonzero elements of R. Choosing an appropriate threshold $\epsilon$ is critical for constructing meaningful recurrence plots, as it determines which state pairs are considered recurrent and thereby influences the actual dynamical state of the system. An optimal threshold preserves significant features of the dynamics while excluding noise-induced or irrelevant recurrences [33]. If $\epsilon$ is too small, the recurrence plot may contain very few recurrence points, leading to the omission of important dynamical patterns. Conversely, if $\epsilon$ is too large, the plot becomes overly dense, which can obscure important structures such as diagonal lines that reflect periodic or quasi-periodic behavior.
@@
To select an appropriate threshold, the common approaches include fixing $\epsilon$ as a percentage of the signal’s standard deviation [2], or adjusting it to maintain a fixed recurrence rate across datasets [31]. In our study, we set the threshold $\epsilon$ to achieve a fixed recurrence rate of $5\%$. To do this, we compute all pairwise distances in the embedded d-dimensional phase space and select the distance value at which $5\%$ of the distances are smaller than or equal to it, as the threshold $\epsilon$.
@@
Figure 2: Representative recurrence plots from all the seven classes. These plots are obtained from the univariate time series corresponding to the following seven classes: (a) Periodic (Lorenz 4D), (b) Quasiperiodic, (c) Chaotic (Lorenz 4D), (d) Hyperchaotic (Chen 4D), (e) Red Noise, (f) White Noise, and (g) Pink Noise.
2.4 Deep Learning Models for Classification
@@
We first implement a single-branch deep learning model based on the ResNet-50 architecture. The model uses a ResNet-50 [16] architecture pre-trained on the ImageNet dataset [43] and is subsequently fine-tuned on RP images. The final feature maps from the backbone are passed through a Global Average Pooling (GAP) layer, followed by a fully connected layer with 1024 neurons, a Batch Normalization layer, and a ReLU activation. A dropout layer (rate = $0.5$) is then applied to reduce overfitting. Finally, a softmax output layer produces the class probability distribution. The model uses categorical cross-entropy loss and the AdamW[29] optimizer, and training is performed with early stopping based on validation accuracy to prevent overfitting.
We also fine-tune a MobileNetV2 architecture pre-trained on ImageNet and subsequently fine-tuned on the RP dataset, using the same transfer-learning setup as ResNet-50. MobileNetV2 exhibits lower performance than ResNet-50, and the classification results produced by both models are not reliable across all data classes.
In this context, we introduce a dual-branch deep learning implementation based on ResNet-50, termed the DBResNet-50 model. The ImageNet-pretrained ResNet-50 backbone is used without its classifier. The standard branch applies global average pooling and $L2$ normalization to the final ResNet-50 features. The directional branch extracts features from the third convolutional layer in stage 2 and an upsampled fourth convolutional layer in stage 3, concatenates them, and processes them with a $1\times 1$ convolution (128 channels, ReLU), a squeeze $1\times 1$ convolution, eight fixed directional $3\times 3$ kernels, batch normalization with ReLU, and three dilated $3\times 3$ convolutions (dilations 1, 2, 3). Its output is globally averaged, batch-normalized, $L2$-normalized, and fed to an auxiliary softmax head. A learnable channel-wise gate scales this branch’s output before concatenation with the standard branch. The fused features pass through a 1024-unit fully connected layer (BN, ReLU, dropout = 0.5) and a final softmax layer.
Training minimizes $L=L_{\text{main}}+0.3L_{\text{aux}}$ using AdamW (learning rate $5\times 10^{-4}$, weight decay $10^{-4}$). Phase 1 trains only newly added layers for 12 epochs; Phase 2 fine tunes the backbone stages from conv3 to conv5 for up to 100 epochs, with early stopping set to stop training if the performance metric does not improve for 10 consecutive epochs. All models are trained with a batch size of 8.
3 Results
The main objective of this work is to accurately identify and classify distinct dynamical behaviors and noisy patterns in data sets, using a machine learning model trained on recurrence plots (RPs). For training, we generate 600 RPs per class from various dynamical systems and noise models, as listed in Table 1. For validation, we generate $100$ RPs per class from the systems in Table 1, ensuring that they are distinct from the training set.
The complete dataset contains seven distinct classes. A few typical representative RPs from each class are presented in Fig. 2, with the images resized to $800\times 800$ pixels.
@@ 3 Results
To enhance prediction accuracy, we apply a test-time augmentation strategy in which each input image is center-cropped by up to $15\%$, while preserving the original aspect ratio. Predictions are generated for all augmented variants of the same image, and the class with the highest average confidence across these predictions is selected as the final output. This ensemble-like approach improves robustness to minor variations in recurrence plot appearance.
@@
To evaluate the performance of the classification models, we compute the confusion matrices and standard classification metrics [52, 40]. In addition, we determine the per-class accuracy, defined as follows: Given a confusion matrix $\mathbf{C}\in\mathbb{N}^{K\times K}$, where $C_{i,j}$ denotes the number of samples whose true class is $i$ and predicted class is $j$, and $K$ denotes the number of classes, the accuracy per class for class $i$ is defined as:
@@
where, $C_{i,i}$ is the number of correctly predicted samples for the class $i$, $\sum_{j=1}^{K}C_{i,j}$ is the total number of samples belonging to the class $i$. This metric reflects how well the model performs for each class individually.
3.1 Classification of simulated time series data
To evaluate the performance of the trained machine learning models, namely, ResNet-50, MobileNetV2, and the proposed DBResNet-50, the models are tested on recurrence plots generated from simulated time series data produced with a different parameter regime that preserves the same class dynamical state as the training dataset. The corresponding performance results for all three models are summarized in Table 2. DBResNet-50 model achieves the highest accuracy of 96.86%, while ResNet-50 achieves 90.14% and MobileNetV2 achieves 86.86%.
Table 2: Performance comparison of models for classifying test data from images of their RPs
Class DBResNet-50 ResNet-50 MobileNetV2
Precision Recall F1-Score Accuracy Precision Recall F1-Score Accuracy Precision Recall F1-Score Accuracy
Chaotic 0.9333 0.8400 0.8842 0.8400 0.8053 0.9100 0.8545 0.9100 0.8571 0.5400 0.6626 0.5400
Hyperchaotic 0.8621 1.0000 0.9259 1.0000 0.9490 0.9300 0.9394 0.9300 0.5839 0.9400 0.7203 0.9400
Periodic 1.0000 0.9800 0.9899 0.9800 0.7143 0.8500 0.7763 0.8500 0.9796 0.9600 0.9697 0.9600
Pink Noise 1.0000 1.0000 1.0000 1.0000 0.9901 1.0000 0.9950 1.0000 0.9252 0.9900 0.9565 0.9900
Quasi-periodic 1.0000 0.9600 0.9796 0.9600 0.9054 0.6700 0.7701 0.6700 1.0000 0.6600 0.7952 0.6600
Red Noise 1.0000 1.0000 1.0000 1.0000 1.0000 0.9500 0.9744 0.9500 0.9709 1.0000 0.9852 1.0000
White Noise 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 0.9706 0.9900 0.9802 0.9900
Accuracy 0.9686 0.9014 0.8686
The confusion matrices in Fig. 3 present an overview of the classification performance across all seven classes for the three models. The diagonal elements indicate correct classifications, while off-diagonal elements represent misclassifications. It is evident that DBResNet-50 is the best-performing model, as nearly all samples lie along the diagonal.
Figure 3: Confusion matrices. Confusion matrices showing classification performance of (a) DBResNet-50 (b) ResNet-50, (c) MobileNetV2 models on datasets generated by simulating models of Table 1 at parameter sets different from training but show same type dynamical behaviour. Diagonal cells indicate correctly classified recurrence plots, while off-diagonal cells indicate misclassifications. It is evident that DBResNet-50 is the best-performing model, as nearly all samples lie along the diagonal.
Table 3: Details of additional dynamical systems used for generating data for validation.
System Equations Parameters Class
Rössler System [17] $\begin{aligned} \dot{x}&=-y-z\\ \dot{y}&=x+ay\\ \dot{z}&=b+z(x-c)\end{aligned}$ a=0.2,b=0.2,c=6\,a=0.2,\;b=0.2,\;c=6\, Chaotic
Duffing System [22] $\begin{aligned} \dot{x}&=y\\ \dot{y}&=x-x^{3}-\delta y+A\sin(\omega t)\end{aligned}$ $\delta=0.5,\;\omega=1,\;A=0.7\,$ Chaotic
Hyperchaotic System [5] $\begin{aligned} \dot{x}&=ax-yz+w\\ \dot{y}&=xz-by\\ \dot{z}&=xy-cz\\ \dot{w}&=-y+d\end{aligned}$ a=8,b=40,c=10,d=0.1\,a=8,\;b=40,\;c=10,\;d=-0.1\, Hyperchaotic
To further assess the robustness and generalization capability of the DBResNet-50 model, we evaluate its performance on recurrence plots generated from dynamical systems not encountered during training. The governing equations and parameter values of these validation systems are summarized in Table 3. For each system, we generate ten time series realizations by varying the initial conditions while keeping the system parameters fixed. The model correctly predicts the underlying dynamics for all test datasets.
We additionally assess the model’s robustness using surrogate data analysis. Shuffled surrogates are generated for all deterministic classes: periodic, quasiperiodic, chaotic, and hyperchaotic using the test datasets listed in Table 1. DBResNet-50 predominantly identifies the surrogate time series as stochastic. This demonstrates that the DBResNet-50 model does not rely on superficial or dataset-specific recurrence plot textures, but is instead sensitive to the disruption of dynamical structure introduced by surrogate randomization.
3.2 Classification of empirical/observational time series data
To further evaluate the generalizability of deep learning models, we test them on experimental or observational time series data. We use interpolation and binning strategies to address missing data and the irregular sampling of real-world time series.
@@ 3.2 Classification of empirical/observational time series data
As a first example, we use time series data obtained experimentally from the Chua circuit. We generate and pass the recurrence plots with three different lengths, $10000,5000$, and $2500$, of time series to the models to test their performance on varying pattern scales within the recurrence plots. The cropping and the corresponding recurrence plots are shown in Fig. 4. The DBResNet-50 model correctly classifies the Chua circuit signal as chaotic for all three cropped ensembles, consistent with the well-established chaotic dynamics of the Chua system [26, 25]. The predictions of the other two models are not correct in all cases as is clear from Table 4.
@@
Figure 4: Experimental data from Chua’s circuit. (a) Time series obtained experimentally from Chua’s circuit. The red lines indicate the cropped portions of the time series at 2,500 and 5,000 data points, while the entire plot corresponds to 10,000 points. (b) Corresponding RP, with the red lines highlighting the regions in the RP corresponding to the cropped portions shown in (a).
Next, we evaluate the models using X-ray light-curve data of the black hole system GRS 1915+105, obtained from the RXTE TOO archive (https://heasarc.gsfc.nasa.gov). The source is reported to exhibit twelve well-known variability classes, each corresponding to a distinct dynamical state characterized by its count-rate evolution and spectral color properties [4]. The data are processed as described in [15]. Among these variability classes, the $\gamma$ and $\chi$ classes are known to exhibit white noise behavior [4, 15]. We select representative data of the light curves from these two classes, and the corresponding light curves and recurrence plots are shown in Fig. 5. The $\chi$ class is correctly classified by all three models, whereas the $\gamma$ class is correctly predicted only by our DBResNet-50 model, as summarized in Table 4.
Figure 5: X-ray light curves from the black hole GRS 1915+105. (a) Light curve data from category $\gamma$ and (b) its corresponding recurrence plot. (c) Light curve data from category $\chi$ and (d) its corresponding recurrence plot.
We further validate the models using the light curve data from three variable stars, AC Her, SX Her, and Chi Cygni, sourced from the AAVSO database (https://www.aavso.org/data-download). Fig. 6 presents the light curves and their corresponding recurrence plots. The light curves for AC Her and SX Her span 14 years, beginning on 20 September 1980, while the Chi Cygni light curve covers a 5-year period starting on 20 September 2019. To ensure consistent sampling and reduce observational noise, we smooth the AC Her and SX Her light curves using a 5-day rolling average, and the Chi Cygni light curve using a 20-day rolling average. All light curves are then interpolated using cubic splines with a smoothing parameter of $s=0.065$. Both the proposed DBResNet-50 model and ResNet-50 correctly classify AC Her and SX Her as chaotic, consistent with earlier studies  [7, 21]. In contrast, MobileNetV2 predicts both stars to be hyperchaotic (cf. Table 4). This suggests that lighter models may struggle to distinguish subtle differences in complex dynamical behavior compared with deeper architectures.
@@
Finally, we evaluate the models using the reanalysis data of temperature from the NCEP (National Centers for Environmental Prediction, https://psl.noaa.gov/), focusing on two locations: Ladakh and Ranchi. The original reanalysis data, sampled four times per day, are preprocessed by binning to two samples per day, followed by removal of the dominant annual cycle. A moving mean with a window size of thirty is applied for smoothing, and the resulting signal is rescaled to the range $[0,1]$.
The DBResNet-50 model classifies both Ranchi and Ladakh as hyperchaotic. However, for Ladakh, the predicted class probabilities exhibit a mixed distribution, with hyperchaos and pink noise contributing approximately $49.7\%$ and $43\%$, respectively. For Ranchi, the model assigns about $90\%$ probability to the hyperchaos class, with all other stochastic categories receiving negligible probabilities. These results are consistent with previously reported findings, which show that for Ladakh, the RQA measures decline sharply before 1980, indicating increasingly stochastic behavior from 1980 to 2000  [18]. In contrast, these measures for Ranchi remain consistently higher over the same period [18].
Thus, DBResNet-50 is shown to demonstrate superior performance in classifying empirical time series data, as compared with the ResNet-50 and MobileNetV2 models. Interestingly, our model is also capable of capturing the mixed nature of the time series, as reflected in its nuanced class-probability distributions.
Table 4: Performance of the deep learning models in classifying observational and experimental datasets.
Time Series True Dynamics DBResNet-50 ResNet-50 MobileNetV2
Chua Circuit 10000 Chaotic [26] Chaotic Periodic Chaotic
Chua Circuit 5000 Chaotic [26] Chaotic Hyperchaotic Red Noise
Chua Circuit 2500 Chaotic [26] Chaotic Chaotic Chaotic
GRS 1915+105 $\chi$ White Noise [4] White Noise White Noise White Noise
GRS 1915+105 $\gamma$ White Noise [4] White Noise Red Noise Red Noise
AC Her Chaotic [21] Chaotic Chaotic Hyperchaotic
Chi Cyg Periodic [48] Periodic Periodic Periodic
SX Her Chaotic [7] Chaotic Chaotic Hyperchaotic
4 Conclusion
@@
In this study, we show that RP images enable classification of seven different dynamical states from data without relying on computed measures using RQA. For this, we introduce a dual branch deep learning model named, DBResNet-50. The architecture of our model builds on a ResNet-50 backbone and incorporates two parallel branches, one that captures global spatial information and another that extracts and emphasizes directional features from early mid-level features, enhancing the model’s classification performance over the standard ResNet-50. The main question we address is one of broad relevance: Can we reliably infer the underlying dynamics of real-world time series data? To achieve this, we generate RPs, extracted via phase-space reconstruction using Takens’ embedding theorem, from nonlinear time series data obtained through simulations of various dynamical systems and noise data. Our proposed DBResNet-50 model takes these RPs as input and learns to identify complex dynamical structures directly from the images of RPs. The model thus classifies these time series into seven distinct classes: periodic, quasi-periodic, chaotic, hyperchaotic, white noise, pink noise, and red noise.
The trained DBResNet-50 model also accurately classifies unseen time series data obtained by simulating dynamical systems for example, Duffing, and Rössler 3D systems, as well as experimental data taken from a Chua circuit, astronomical light curves from the black-hole GRS 1915+105, AC Her, SX Her, and Chi Cygni, as shown in Table 4. The models also provide insight into the amount of stochasticity in the data through the class wise probabilities, helping us to interpret mixed or ambiguous regimes in the climate data as demonstrated in the case of temperature time series from Ladakh and Ranchi. We also explore the sensitivity of the empirical results by evaluating the model, trained using a 5% recurrence rate (RR) on the same empirical data while generating recurrence plots (RPs) with 3% and 7% RR. We do not observe any considerable differences in the results, implying that small variations in the recurrence rate can also be accommodated by the model.
In summary, our framework demonstrates robust classification performance across synthetic, experimental, and observational datasets. This demonstrates the model’s applicability to real-world data and reinforces the potential of using recurrence based deep learning models for determining the inherent dynamics of the system. We also compare the performance of proposed DBResNet-50 implementation with other deep learning models namely, ResNet-50, and MobileNetV2. We find that the proposed DBResNet-50 model performs best, followed by ResNet-50, while MobileNetV2 shows comparatively lower accuracy.
@@
Compared with traditional approaches that rely on recurrence quantification measures and recurrence network metrics [50], our method offers a more efficient alternative. It eliminates the need for computing recurrence measures and feature selection by learning discriminative features directly from RP images. Our results show that high classification accuracy can be achieved without computing various recurrence measures.
@@
Looking ahead, several directions can further extend the capabilities of our model. One potential approach is to explore an ensemble RR method in which the model processes recurrence plots generated with multiple recurrence rates to make a single prediction, helping bridge the gap across studies that use different RR approaches. Another interesting application involves detecting dynamical transitions within time series by segmenting recurrence plots along the line of identity (LOI) using our model on transition-rich data to identify these substructures, similar to time-dependent RQA techniques [33]. Another enhancement includes combining our approach with model-based generative frameworks, such as delay embedding and deep delay autoencoders [3], to synthesize training data from partial observations of unknown systems. This could reduce dependence on fully simulated datasets and improve performance under limited data conditions.
In conclusion, the reported study establishes and explores deep learning approaches that classify seven distinct dynamical and noise regimes directly from recurrence plots, providing a powerful, efficient, and generalizable framework for identifying dynamical states in synthetic, astronomical, climate, and electronic circuit time series data.
5 Acknowledgments